OpenStudy (anonymous):

Parker earned $32 by selling 27 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold.

OpenStudy (anonymous):

I dont think anyone is just gonna give you the answer

OpenStudy (anonymous):

Actually, I didn't want an answer. I need HELP, Not trying to cheat. But thanks for your unwanted imput. :)

OpenStudy (anonymous):

Might wanna edit and erase "i just need the answer", then.

OpenStudy (anonymous):

Look, if you're not going to help * me, then bye.

OpenStudy (anonymous):

A model would be an equation that includes the info from the problem... Here are the key bits of information the problem provides: $32 total 27 glasses sold (includes large and small glasses) $1.50 per large glass $1 per small glass So, just to get an idea of how to set this up, think of how much he makes if he sells one large and one small glass... total = (1 x $1) + (1 x $1.50) = $2.50 Or 2 large glasses and three small glasses: total = (3 x $1) + (2 x $1.50) = $6.00 So, $32 = L x $1.50 + S x $1 where L is the number of large glasses and S is the number of small glasses And you also know that the total number of glasses sold is 27. So L + S = 27 That means your model is these 2 equations: 1.50L + S = 32 L + S = 27

OpenStudy (anonymous):

thank youuuu.

OpenStudy (anonymous):

Now that you have those 2 equations, you somehow have to find a way to get the number of large glasses sold, since that's what the problem asks... So, you might notice that in the 2nd equation, if L + S = 27, you could rewrite it as S = 27 - L. If the total number of glasses sold is 27, then the number of small glasses, S, is 27 minus the number of large glasses, L. You still don't know what L is, but that's ok... Since you have a "nickname" for S... you can replace any S with "27 - L"... then use that in the other equation: 1.50L + S = 32 1.50L + (27-L) = 32 <<-- replaced "S" with "27 - L" 1.5L + 27 - L = 32 1.5L - L + 27 - 27 = 32 - 27 <<<--- subtracted 27 from both sides 0.5L = 5 <<<---- simplified both sides.... 2(0.5L) = 2(5) <<<--- multiplied both sides by 2 L = 10

OpenStudy (anonymous):

Thanks so much, You're so amazing. I understand everything now! (:

OpenStudy (anonymous):

It's a bunch of steps, but read through them... each one is pretty simple, and step by step, you work through until you have solved for L, the number of large glasses sold.

OpenStudy (anonymous):

Thank you! (:

OpenStudy (anonymous):

Glad to help... hope you learned the idea a little :) It takes practice, but you can get these :) Good luck...

OpenStudy (anonymous):

I get it like perfectly. And I read everything. You're so great!(:

OpenStudy (anonymous):

Thanks -- I know it's a lot to look at, but I believe it's better to help you learn if you see the process than if you just get a quick answer. Keep it up :)